# Non-Maxwellianity of Ion Velocity Distributions in the Earth’s Magnetosheath

### v1

@@ Abstract
This classification reveals two distinct clusters at $(\beta_{i}\sim 1,~{}B_{\mathrm{rms}}/B_{0}\sim 0.1)$ and $(\beta_{i}\sim 10,~{}B_{\mathrm{rms}}/B_{0}\sim 1)$, indicating a strong correlation between turbulence levels and ion plasma beta. These clusters correspond to the quasi-perpendicular and quasi-parallel magnetosheath, respectively, as demonstrated by Ref. [29]. We find that the temperature anisotropy, $T_{i\bot}/T_{i\parallel}$, decreases with increasing $\beta_{i}$ and $B_{\mathrm{rms}}/B_{0}$ [Fig. 3(b)]. However, due to the correlation between $\beta_{i}$ and $B_{\mathrm{rms}}/B_{0}$, and the lack of data at ($\beta_{i}\sim 1,~{}B_{\mathrm{rms}}/B_{0}\sim 1$) and ($\beta_{i}\sim 10,~{}B_{\mathrm{rms}}/B_{0}\sim 0.1$), it remains unclear whether this decrease is a direct effect of turbulence or an indirect consequence of its relationship with $\beta_{i}$.
In contrast, the agyrotropy $\sqrt{Q_{i}}$ [Fig. 3(c)] and the non-bi-Maxwellianity $\varepsilon_{i}$ decrease with $\beta_{i}$ but increase with $B_{\mathrm{rms}}/B_{0}$ [Fig. 3(c), 3(d)]. Notably, at $\beta_{i}\sim 5,~{}B_{\mathrm{rms}}/B_{0}\sim 1$, we find values reaching $\sqrt{Q_{i}}\simeq 0.07$ and $\varepsilon_{i}\simeq 0.5$, which exceed the dataset mean by one standard deviation ($\sqrt{Q_{i}}=0.04^{+0.02}_{-0.02}$, $\varepsilon_{i}=0.36^{+0.11}_{-0.07}$) [Figs. 1(b), 1(c)]. Furthermore, the eigenvalue ratios $\lambda_{1}/\lambda_{2}$ and $\lambda_{1}/\lambda_{3}$ closely follow the trends of agyrotropy and anisotropy, respectively [Figs. 3(b), 3(c), 3(e), 3(f)]. This indicates that in the low-$\beta_{i}$, weakly turbulent regime, iVDFs are compressed along the magnetic field, i.e., $\lambda_{1}/\lambda_{3}>1$, with little high-order non-Maxwellianity. In contrast, in the high-$\beta_{i}$, strongly turbulent regime, the iVDFs become nearly isotropic with pronounced high-order non-Maxwellian features.
To understand how iVDFs relax toward LTE, we examine the magnetic field curvature relative to the ion scale. In addition to wave-particle interactions, ion scattering in strongly curved magnetic fields can efficiently isotropize iVDFs [40, 22, 5]. We compute the magnetic field curvature $\bm{K}=\hat{\bm{b}}\cdot\nabla\hat{\bm{b}}$ using the multi-spacecraft curlometer technique [30]. The PDF of the normalized magnetic curvature $|\bm{K}|/|\bm{K}|_{\mathrm{rms}}$ follows a double Pareto lognormal distribution [Fig. 4a]. Notably, the distribution peaks at $|\bm{K}|^{\star}\simeq 0.14|\bm{K}|_{\mathrm{rms}}\simeq 0.02d_{i}^{-1}$ (red circle), which corresponds to $|\bm{K}|^{\star}\sim\lambda_{c}^{-1}$, with $\lambda_{c}\sim 10-100d_{i}$ [28]. Moreover, for $|\bm{K}|\ll|\bm{K}|^{\star}$ and $|\bm{K}|\gg|\bm{K}|^{\star}$, the distribution follows power-law slopes of 1 and $-2.5$, respectively, confirming theoretical predictions and earlier findings [41, 42].
@@
We compute the adiabaticity parameter, $\kappa=\sqrt{r_{c}/\rho_{i}}$, where $r_{c}=1/|\bm{K}|$ is the radius of curvature of the magnetic field, as a proxy for ion motion dynamics [40] [Fig. 4b]. For $\kappa\gg 1$, ion motion is adiabatic, with the magnetic moment $\mu$ conserved to an accuracy of $\Delta\mu/\mu\sim\exp(-2\kappa^{2}/3)$ [43, 40]. At $\kappa\sim 1$, resonance occurs, leading to chaotic ion motion and strong scattering, characterized by $\Delta\mu/\mu\sim 1$ [43, 40]. For $\kappa\ll 1$, the motion is quasi-adiabatic, with the generalized magnetic moment $I_{z}$ approximately conserved as $\Delta I_{z}/I_{z}\sim\kappa$ for compressional and $\sim 1$ for force-free discontinuities [44, 45]. We find that for 25% of the iVDFs, the Larmor radius satisfies $\kappa\leq 1.8$ (red shaded area), corresponding to $\Delta\mu/\mu\geq 0.1$ (red circle). Since $\kappa$ is estimated using the thermal ion speed $v_{thi\bot}$, ions with $v_{i\bot}>v_{thi\bot}$ may remain non-adiabatic even when $\kappa\gg 1$. This indicates that a substantial fraction of the ions follow quasi-adiabatic and chaotic orbits, breaking the adiabatic invariant conservation.
@@
The adiabaticity parameter $\kappa$ decreases with increasing $\beta_{i}$ and $B_{\mathrm{rms}}/B_{0}$ due to its dependence on ion temperature and magnetic field strength [Fig. 4c]. Notably, its variation as a function of $\beta_{i}$ and $B_{\mathrm{rms}}/B_{0}$ closely follows the trend of temperature anisotropy [Figs. 3(b), 3(f)]. Specifically, at $(\beta_{i}\sim 1,~{}B_{\mathrm{rms}}/B_{0}\sim 0.1)$, we find $\kappa\simeq 7$, corresponding to $\Delta\mu/\mu\simeq 2\times 10^{-16}$, indicating negligible scattering due to magnetic field curvature. In contrast, at $(\beta_{i}\sim 100,~{}B_{\mathrm{rms}}/B_{0}\sim 1)$, $\kappa\sim 1$, leading to $\Delta\mu/\mu\simeq 1$, and ions experience strong scattering. Moreover, $\kappa$ is strongly correlated with the eigenvalue ratio $\lambda_{1}/\lambda_{3}$ [Fig. 4d], with its 95th percentile reaching a minimum at $\kappa\simeq 1.21$, corresponding to $\Delta\mu/\mu\simeq 0.38$, near the expected maximum scattering at $\kappa=1$. This suggests that the isotropization of iVDFs results from the breakdown of adiabatic invariance due to magnetic field curvature of the order of the ion gyroradius.
@@
Discussion - Our analysis shows that iVDFs in Earth’s magnetosheath can deviate significantly from Maxwell-Boltzmann LTE. Many iVDFs are predicted to be unstable to temperature anisotropy-driven instabilities. However, both stable and unstable iVDFs also display notable high-order non-bi-Maxwellian features, suggesting that the bi-Maxwellian model commonly used in linear Vlasov theory may be insufficient to assess the iVDFs’ stability [18, 19, 20]. Non-LTE iVDFs are often associated with large amplitude magnetic field fluctuations. Interestingly, these fluctuations may be ion scale magnetic shears, pushing the iVDFs further from LTE rather than relaxing them toward marginal stability, as one might expect from electromagnetic waves [46, 39, 47, 48]. Alternatively, these unstable iVDFs may originate at the bow shock and persist across the magnetosheath due to the plasma’s slow transit time across the magnetosheath compared with the relaxation timescales [5].
@@
We demonstrate that the non-Maxwellianity of iVDFs depends on ion plasma beta $\beta_{i}$ and turbulence level $B_{\mathrm{rms}}/B_{0}$. As $B_{\mathrm{rms}}/B_{0}$ increases, high-order non-Maxwellian features, including agyrotropy, become more pronounced. This suggests that turbulence drives these higher-order features. Numerical simulations and in situ observations support the idea that turbulence-generated current sheets play a key role in creating such features [7, 6, 8]. Additionally, the distribution of agyrotropy follows a non-Gaussian power-law tail, which mirrors the behavior of magnetic field increments in turbulence [49, 50]. This further indicates that turbulence is responsible for driving these high-order non-LTE features. Moreover, the fine-scale velocity-space features we measure may be tied to a concurrent phase-space cascade [51], further highlighting the complex interplay between turbulence and iVDF evolution.
@@
Interestingly, despite these high-order non-Maxwellian features, iVDFs in weakly magnetized, highly turbulent regime $(\beta_{i}\gg 1,~{}B_{\mathrm{rms}}/B_{0}\sim 1)$ tend to be nearly isotropic. In these conditions, magnetic curvature induced by turbulence can be of the order of the thermal ion gyroradius scale or smaller. As a result, ions interact with these turbulence-generated curved magnetic fields, following chaotic orbits and undergoing pitch-angle scattering [43, 40, 52, 22, 53]. Numerical simulations further suggest that such strongly curved magnetic fields enhance particle transport and diffusion [54, 55, 56]. This indicates that turbulence distorts magnetic fields across all scales down to sub-ion scales, leading to phase-space diffusion and ion energization as ions interact with these curved fields.
Conclusion - We present the first statistical quantitative analysis of the non-Maxwellianity of iVDFs in the turbulent collisionless plasma in the Earth’s magnetosheath. Our results indicate that the shocked plasma rapidly relaxes. We show that the complexity of non-LTE iVDFs depends on the ion plasma beta and turbulence intensity. As magnetic field fluctuations increase, iVDFs exhibit substantial high-order deviations from LTE, while wave excitation and curvature scattering limit low-order non-LTE. These results highlight the role of turbulence in plasma heating and driving and regulating non-LTE behaviors in iVDFs, providing insights into its interaction with ion velocity distributions in collisionless plasmas.
@@ Acknowledgements.
Acknowledgments - We thank the MMS team for data access and support. L.R. acknowledges support from the Knut and Alice Wallenberg Foundation (Dnr. 2022.0087), the Swedish National Space Agency grant 192/20, and the Royal Swedish Academy of Sciences grant AST2024-0015. S.S. acknowledges the Space It Up project funded by the Italian Space Agency, ASI, and the Ministry of University and Research, MUR, under contract n. 2024-5-E.0 - CUP n. I53D24000060005. I.S. acknowledges support from the Swedish Research Council Grant 2016‐0550 and the Swedish National Space Agency Grant 158/16. A.C. acknowledges support from NASA Grants 80NSSC21K0454, 80NSSC22K0688, and 80NSSC24K0172. O.P. acknowledges the project ”2022KL38BK - The ULtimate fate of TuRbulence from space to laboratory plAsmas (ULTRA)” (Master CUP B53D23004850006), by the Italian Ministry of University and Research, funded under the National Recovery and Resilience Plan (NRRP), Mission 4 – Component C2 – Investment 1.1, ”Fondo per il Programma Nazionale di Ricerca e Progetti di Rilevante Interesse Nazionale (PRIN 2022)” (PE9) by the European Union – NextGenerationEU. This research was supported by the International Space Science Institute (ISSI) in Bern through the ISSI International Team project #23-588 (”Unveiling Energy Conversion and Dissipation in Non-Equilibrium Space Plasmas”). L.R. thanks Yu. V. Khotyaintsev for helpful suggestions. Data analysis used the pyrfu analysis package  [^1] See https://pypi.org/project/pyrfu/..

### v2

@@ Abstract
This classification reveals two distinct clusters at $(\beta_{i}\sim 1,~{}B_{\mathrm{rms}}/B_{0}\sim 0.1)$ and $(\beta_{i}\sim 10,~{}B_{\mathrm{rms}}/B_{0}\sim 1)$, corresponding to the quasi-perpendicular and quasi-parallel magnetosheath, respectively [32]. We find that the temperature anisotropy, $T_{i\bot}/T_{i\parallel}$, decreases with increasing $\beta_{i}$ and $B_{\mathrm{rms}}/B_{0}$ [Fig. 3(b)]. Due to the correlation between $\beta_{i}$ and $B_{\mathrm{rms}}/B_{0}$, and the lack of data at ($\beta_{i}\sim 1,~{}B_{\mathrm{rms}}/B_{0}\sim 1$) and ($\beta_{i}\sim 10,~{}B_{\mathrm{rms}}/B_{0}\sim 0.1$), it remains unclear whether this decrease is a direct effect of turbulence or an indirect consequence of its relationship with $\beta_{i}$.
In contrast, the agyrotropy $\sqrt{Q_{i}}$ [Fig. 3(c)] and the non-bi-Maxwellianity $\varepsilon_{i}$ decrease with increasing $\beta_{i}$ but increase with $B_{\mathrm{rms}}/B_{0}$ [Fig. 3(c), 3(d)]. Notably, at $\beta_{i}\sim 5,~{}B_{\mathrm{rms}}/B_{0}\sim 1$, we find values reaching $\sqrt{Q_{i}}\simeq 0.07$ and $\varepsilon_{i}\simeq 0.5$, which significantly exceed the dataset mean ($\sqrt{Q_{i}}=0.04^{+0.02}_{-0.02}$, $\varepsilon_{i}=0.36^{+0.11}_{-0.07}$) [Figs. 1(b), 1(c)]. Furthermore, the eigenvalue ratios $\lambda_{1}/\lambda_{2}$ and $\lambda_{1}/\lambda_{3}$ closely follow the trends of agyrotropy and anisotropy, respectively [Figs. 3(b), 3(c), 3(e), 3(f)]. This indicates that in the low-$\beta_{i}$, weakly turbulent regime, iVDFs are compressed along the magnetic field, i.e., $\lambda_{1}\sim\lambda_{2}>\lambda_{3}$, with little high-order non-bi-Maxwellianity. In contrast, in the high-$\beta_{i}$, strongly turbulent regime, the iVDFs become globally isotropic with pronounced high-order non-bi-Maxwellian features.
To understand how iVDFs relax toward LTE, we examine the magnetic field curvature relative to the ion gyroradius. Indeed, in addition to wave-particle interactions, sharp magnetic field bends can efficiently scatter ions [47, 27, 5, 28]. The PDF of the normalized magnetic curvature $|\bm{K}|/|\bm{K}|_{\mathrm{rms}}$ resembles a double Pareto lognormal distribution [Fig. 4a], supporting numerical simulations and earlier observations [48, 49]. Notably, the PDF peaks at $|\bm{K}|^{\star}\simeq 0.14|\bm{K}|_{\mathrm{rms}}\simeq 0.02d_{i}^{-1}$ (red circle), which corresponds to $|\bm{K}|^{\star}\sim\lambda_{c}^{-1}$, with $\lambda_{c}\sim 10-100d_{i}$ [36]. Moreover, for $|\bm{K}|\ll|\bm{K}|^{\star}$ and $|\bm{K}|\gg|\bm{K}|^{\star}$, the PDF approximately follows power-law slopes of $1$ and $-2.5$, respectively. Theoretical predictions suggest that these asymptotic limits result from the large curvature $\bm{K}$ where the magnetic field strength is weak, and the low curvature associated with small magnetic tension forces, respectively [48, 49].
@@
We compute the adiabaticity parameter, $\kappa=\sqrt{r_{c}/\rho_{i}}$, where $r_{c}=1/|\bm{K}|$ is the radius of curvature of the magnetic field, as a proxy for ion motion dynamics [47] [Fig. 4b]. For $\kappa\gg 1$, ion motion is adiabatic, with the magnetic moment $\mu$ conserved to an accuracy of $\Delta\mu/\mu\sim\exp(-2\kappa^{2}/3)$ [50, 47]. At $\kappa\sim 1$, resonance occurs, leading to chaotic ion motion and strong scattering, characterized by $\Delta\mu/\mu\sim 1$ [50, 47]. For $\kappa\ll 1$, the motion is quasi-adiabatic, with the generalized magnetic moment $I_{z}$ approximately conserved as $\Delta I_{z}/I_{z}\sim\kappa$ for compressional and $\sim 1$ for force-free discontinuities [51, 52]. We find that for 25% of the iVDFs $\kappa\leq 1.8$ (red shaded area), corresponding to $\Delta\mu/\mu\geq 0.1$ (red circle). Since $\kappa$ is estimated using the thermal ion speed $v_{thi\bot}$, ions with $v_{i\bot}>v_{thi\bot}$ may remain non-adiabatic even when $\kappa\gg 1$. This indicates that a substantial fraction of the ions follow non-adiabatic orbits, thereby breaking the conservation of magnetic moment.
@@
The adiabaticity parameter $\kappa$ decreases with increasing $\beta_{i}$ and $B_{\mathrm{rms}}/B_{0}$ due to its temperature and magnetic field strength dependence [Fig. 4c]. Notably, its variation with $\beta_{i}$ and $B_{\mathrm{rms}}/B_{0}$ closely follows the trend of $T_{i\bot}/T_{i\parallel}$ [Figs. 3(b), 3(f)]. At $(\beta_{i}\sim 1,~{}B_{\mathrm{rms}}/B_{0}\sim 0.1)$, we find $\kappa\simeq 7$, corresponding to $\Delta\mu/\mu\simeq 2\times 10^{-16}$, indicating negligible curvature scattering. In contrast, at $(\beta_{i}\sim 100,~{}B_{\mathrm{rms}}/B_{0}\sim 1)$, $\kappa\sim 1$, leading to $\Delta\mu/\mu\simeq 1$, and ions experience strong scattering. Moreover, $\kappa$ is strongly correlated with $\lambda_{1}/\lambda_{3}$ [Fig. 4d], whose 95th percentile reaches a minimum at $\kappa\simeq 1.21$, corresponding to $\Delta\mu/\mu\simeq 0.38$, near the expected maximum scattering at $\kappa=1$. This suggests that the isotropization of iVDFs results from the violation of adiabatic invariance due to sharp magnetic field bends.
@@
Discussion - Our analysis shows that iVDFs in Earth’s magnetosheath can deviate significantly from Maxwell-Boltzmann LTE. Many iVDFs are predicted to be unstable to temperature anisotropy-driven instabilities. However, both stable and unstable iVDFs also display notable high-order non-bi-Maxwellian features, suggesting that the bi-Maxwellian model commonly used in linear Vlasov theory may be insufficient to assess the iVDFs’ stability [53, 54, 55, 20]. We note that we do not account for contributions from other species, such as alpha particles, which may influence the non-Maxwellianity and the stability limits [56, 57, 54, 58]. Non-LTE iVDFs are often associated with large amplitude magnetic field fluctuations. These fluctuations may be ion scale magnetic shears, pushing the iVDFs further from LTE rather than relaxing them toward marginal stability, as one might expect from electromagnetic waves [59, 46, 60, 61]. Alternatively, these unstable iVDFs may originate at the bow shock and persist due to the plasma’s slow transit time across the magnetosheath compared with the relaxation timescales [5].
@@
We demonstrate that the non-Maxwellianity of iVDFs depends on $\beta_{i}$ and the turbulence level $B_{\mathrm{rms}}/B_{0}$. As $B_{\mathrm{rms}}/B_{0}$ increases, high-order non-Maxwellian features, including agyrotropy, become more pronounced, suggesting that these are driven by the turbulence. Numerical simulations and in situ observations suggest that turbulence-generated current sheets create such features [7, 6, 8]. Additionally, the agyrotropy follows a non-Gaussian power-law distribution, mirroring the behavior of magnetic field increments in turbulence [62, 63]. This suggests turbulence may be linked to these high-order non-LTE features. Moreover, the measured fine-scale velocity-space features may be tied to a concurrent phase-space cascade [1, 64, 65], further highlighting the complex interplay between turbulence and iVDF evolution.
@@
Interestingly, in weakly magnetized, highly turbulent regime $(\beta_{i}\gg 1,~{}B_{\mathrm{rms}}/B_{0}\sim 1)$ iVDFs are globally isotropic despite high-order non-Maxwellian features. We note that while the iVDF is globally isotropic, it may contain anisotropic parts balancing each other. Nevertheless, in these conditions, magnetic curvature induced by turbulence can be of the order of the ion gyroradius scale or smaller. As a result, ions interact with the turbulence-generated curved magnetic fields, undergoing pitch-angle scattering [50, 47, 66, 27, 67]. Numerical simulations further suggest that such strongly curved magnetic fields enhance particle transport and diffusion [68, 69, 28]. This indicates that turbulence distorts magnetic fields across all scales down to sub-ion scales, leading to ion phase-space diffusion and energization through interaction with these curved fields.
Conclusion - We present the first quantitative, statistical analysis of iVDFs in the turbulent, collisionless plasma of Earth’s magnetosheath. Using a framework to quantify deviations from LTE, we show that departures from a bi-Maxwellian are both widespread and significant. Our novel parametric analysis reveals that these deviations decrease with $\beta_{i}$ and increase with the turbulence level. The ubiquity of such non-Maxwellian features raises a question about the degree to which the bi-Maxwellian-based theory is applicable in describing the relaxation of the shocked turbulent plasma.
We demonstrate that curvature scattering in the presence of large-amplitude turbulence contributes to the relaxation of non-LTE iVDFs, pointing to alternative, turbulence-driven relaxation pathways. These findings advance our understanding of the kinetic response of ions to turbulence and highlight the importance of turbulence-driven processes in shaping the velocity space dynamics of space and astrophysical plasmas.
@@ Acknowledgements.
Acknowledgments - We thank the MMS team for data access and support. L.R. thanks D. J. Gershman for helpful suggestions. L.R. acknowledges support from the Knut and Alice Wallenberg Foundation (Dnr. 2022.0087), the Swedish National Space Agency grant 192/20, and the Royal Swedish Academy of Sciences grant AST2024-0015. S.S. acknowledges the Space It Up project funded by the Italian Space Agency, ASI, and the Ministry of University and Research, MUR, under contract n. 2024-5-E.0 - CUP n. I53D24000060005. I.S. acknowledges support from the Swedish Research Council Grant 2016‐0550 and the Swedish National Space Agency Grant 158/16. K. G. K. was supported by NASA ECIP Grant No. 80NSSC19K0912. A.C. acknowledges support from NASA Grants 80NSSC21K0454, 80NSSC22K0688, and 80NSSC24K0172. O.P. acknowledges the project ”2022KL38BK - The ULtimate fate of TuRbulence from space to laboratory plAsmas (ULTRA)” (Master CUP B53D23004850006), by the Italian Ministry of University and Research, funded under the National Recovery and Resilience Plan (NRRP), Mission 4 – Component C2 – Investment 1.1, ”Fondo per il Programma Nazionale di Ricerca e Progetti di Rilevante Interesse Nazionale (PRIN 2022)” (PE9) by the European Union – NextGenerationEU. This research was supported by the International Space Science Institute (ISSI) in Bern through the ISSI International Team project #23-588 (”Unveiling Energy Conversion and Dissipation in Non-Equilibrium Space Plasmas”). Data analysis used the pyrfu analysis package  [^1] See https://pypi.org/project/pyrfu/..